y =2x^2-12x-20 Is The Equation, And I Don't Understand How To Find The Answer
Answer by fcabanski(1391)

Check my profile and contact me for one on one online tutoring.


The goal is to find the roots of the equation. That is, the value of x when y=0:


0 = 2x^2-12x-20 (the first thing we can do is factor out 2.)


0=x^2-6x-10 (We cannot easily factor this with reverse foil, so we need to use the quadratic equation (-b + or - sqrt (b^2-4ac))/2a:


a=1, b=-6, c=-10.


(-(-6) + or - sqrt (-6^2 - 4(1)(-10)))/2(1)=


(6 + or - sqrt (36+40))/2 =


(6 + or - sqrt(76))/2 =


(6 + or - sqrt(4*19))/2 =


(6 + or - 2sqrt(19))/2 =


3 + or - sqrt(19)